Perpendicular to y=-1/2x+3, passing through the point (3,-4)
Found 3 solutions by mananth, jim_thompson5910, stanbon:
Answer by mananth(16949)

1 y = 0.5 x + 3
Divide by 1
y = 0.5 x + 3
Compare this equation with y=mx+b
slope m = 0.5

The slope of a line perpendicular to the above line will be the negative reciprocal -2
m1*m2=-1
The slope of the required line will be -2.00

m= -2 ,point ( 3 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = -6.00 + b
b= 2
m= -2
The required equation is y = -2 x + 2
m.ananth@hotmail.ca

Answer by jim_thompson5910(35256)

Original slope = -1/2
Perpendicular slope = 2 (negative reciprocal of original slope)


y = mx+b

y = 2x+b

-4 = 2(3)+b

-4 = 6+b

-4-6 = b

-10 = b

So the answer is y = 2x-10

Answer by stanbon(75887)

Perpendicular to y=-1/2x+3, passing through the point (3,-4)
----
That line has a slope of -1/2
---
Any perpendicular line must have slope = 2
---
Find equation of line with m = 2 and y = -4 when x = 3
-----
Form: y = mx + b
Solve for "b":
-4 = 2*3 + b
b = -10
----
Equation:
y = 2x -10
===============
Cheers,
Stan H.
-------------------